• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 00:04:06 on October30,2002

    Entry number 89281

    keyword=swing shift summary

    Run List:
    started 2845

    17:05 MCC will resume low current cw, hope to be back soon.

    17:20 high current (99uA) has returned, production has started.

    19:35 coming back up after ep measurement. MCC calls and says
    that G0 doesn't want us back on line until after they finish some scan (BCM?)
    in their hall, 10-15 minutes. Poetic justice... some power supply in arc 1 just
    crapped out, and no one is getting any beam. Ha! MCC estimates 30 minutes.

    20:30 Beam still not back, MCC estimates another 30 minutes.
    21:15 We're getting closer. MCC now estimates only 20 minutes!
    22:20 Beam has returned!
    22:48 MCC calls to say they are taking away beam for 2 minutes. They are having
    a problem with the Hall B beam, and wonder if its related to our beam.
    22:50 MCC calls to ask for 5 more minutes (in addition to the previous 2, I confirmed) in order to allow Hall B to do complete a harp scan. They see a tail,
    which is suspected to come from bleedthrough from the Hall A beam.
    22:55 MCC was honest!!! Beam returns.
    23:40 Beam has been quite unstable... MCC says trips in the Hall C line is tripping beam in all halls.
    23:59 MCC calls. Beam gone for ~15 minutes (problem with "trim rack", but I didn't catch any other specifics).